6/26/14: After acid wash redeployed Montara active sampler.

6/27/14: Sample bottle found on ground, slipped out of sample tube. System checked and rinsed and new sample bottle deployed.

6/28/14: 100 mL sample collected, preserved and refrigerated. System was still  quite clean after just one day so we just removed and cleaned the racks in DI water  and gave everything a thorough DI rinse. Collected a blank, reset system.

140624 UCSC Observations

Jeremiah arrived at the ISB Greenhouse at 8:45 am.  There was no fog sample collected and the collector doors had opened.  He removed the strand frames, teflon tray, teflon side pieces, top plate, and drip tube assembly.  He washed them with 409 and water and gave a DI rinse.  He reassembled and took a blank sample (B-UCSC-140624).  The system was reset and ready to go for sampling tonight.

140624 LML Observations

Kona and Peter met at the LML site at 10:30 am.  No fog was collected, although the doors had opened.  We removed the strand frames and washed with DI at the sink.  We rinsed the collector body and drip tray/tube assembly in place with copious amounts of DI water.  Reassembled and reset for sampling tonight.